OpenStudy (anonymous):

Simplify using only positive exponents. (-3x^-6)(7x^-7)= -21/x^13

OpenStudy (anonymous):

\[\Large (-3x^{-6})(7x^{-7})=\frac{-21}{x^{13}}\] use this rule... \[\LARGE x^a\cdot x^b=x^{a+b}\] ...multiply left side... you'll get a number, than use cross multiplication and there you go... !